# Implementing Agents in JavaScript

### v1

@@
Abstract
This chapter gives an introduction to agent-oriented programming in JavaScript. It provides an example-based walk-through of how to implement abstractions for reasoning loop agents in vanilla JavaScript. The initial example is used as a stepping stone for explaining how to implement slightly more advanced agents and multi-agent systems using JS-son, a JavaScript library for agent-oriented programming. In this context, the chapter also explains how to integrate reasoning loop agents with generative AI technologies—specifically, large language models. Finally, application scenarios in several technology ecosystems and future research directions are sketched.
@@ 4.1 A Simple Agent in Vanilla JavaScript
First, we model the beliefs of the agent. {minted}[fontsize=]js const beliefs = door: locked: true , requests: [] As we can see, the initial state of the belief base models that the door is currently locked, and that no requests are to be handled. Let us now specify the agent’s plans: i) if the door is not locked and a request to lock the door exists (head) then the agent should act to lock the door (body); ii) if the door is locked and a request to unlock the door exists (head) then the agent should act to unlock the door (body). {minted}[fontsize=]js const plans = [ head: beliefs =¿ !beliefs.door.locked & beliefs.requests.includes(’lock’), body: () =¿ ’lock’ , head: beliefs =¿ beliefs.door.locked & beliefs.requests.includes(’unlock’), body: () =¿ ’unlock’ ] With beliefs and plans specified, we can move on to implement the agent’s reasoning loop. We start by implementing the belief revision function, which updates the agent’s belief about the state of the door, removes requests that have been successfully executed, and adds new requests that have come in. {minted}[fontsize=]js const revise = (beliefs, percepts) =¿ return door: percepts.door, requests: beliefs.requests.filter( request =¿ !percepts.executions.includes(request) ).concat(percepts.requests) Next, we implement the plan deliberation function. The function first determines which bodies of the plans should be activated (deliberate) to then execute these bodies (act). {minted}[fontsize=]js const deliberate = (beliefs, plans) =¿ return activatedBodies = plans.filter( plan =¿ plan.head(beliefs) ).map(plan =¿ plan.body(beliefs)) Finally, we can complete the agent, by i) assigning beliefs and plans, as well as the revision and deliberation functions to an agent object, and ii) implementing a run function that triggers first belief revision and then plan deliberation. {minted}[fontsize=]js const agent = beliefs, plans, revise, deliberate, get run() this.beliefs = this.revise(this.beliefs, percepts) return this.deliberate( this.beliefs, this.plans ) We can observe that we need to make use of a get function to access object-level properties and that percepts will need to be specified outside of the scope of the agent object or any of its functions. This indicates that providing generic abstractions in the form of a library may be useful.
@@
We still need to implement the environment that provides our agent with percepts and processes its actions. {minted}[fontsize=]js const environment = agent, state: door: locked: true , requests: [] , get run() […Array(steps).keys()].forEach( step =¿ this.state.requests = Math.random() ¡ 0.5 ? [’lock’] : [’unlock’] percepts = this.state const exec = this.agent.run const actions = exec ? exec : [] if (actions.includes(’lock’)) this.state.door.locked = true if (actions.includes(’unlock’)) this.state.door.locked = false this.state.executions = actions ) Our environment processes all agent actions by changing the state of the door accordingly and adds pseudo-random requests to either lock or unlock the door, independently of the door’s current status. In this way, the environment primitively simulates the presence of additional agents. Note that in this simple example, our environment is a single agent environment, i.e., it is not programmed to handle more than one agent. We again exploit the get syntax in a way that is arguably not idiomatic.
@@
We can now execute the environment, first specifying the number of environment-agent interactions (steps) it should run. {minted}[fontsize=]js const steps = 20 environment.run When running the example, we see that our porter agent processes the requests it receives from the environment as intended, generating the desired actions, which are then further processed accordingly by the environment [^5] Cf. Appendix Item 3.. When debugging nuances of the example, we observe that nonsensical requests (the locking of locked doors and the unlocking of unlocked doors) “pile up” until they can be meaningfully addressed.
@@ 4.2 An LLM Agent Based on Reusable Abstractions
We first implement the agent’s beliefs, by instantiating JS-son Belief objects. {minted}[fontsize=]js const beliefs = …Belief(’rejectExps’, []), …Belief(’excuseAccepted’, false), …Belief(’name’, ’Bart’), …Belief(’teacherName’, ’Edna Krabappel’) In our case, the result is roughly the same as when specifying beliefs as JavaScript objects. However, in advanced scenarios, JS-son beliefs have in-built support for specifying revision behavior, such as priorities and inter-belief dependencies, where the value of one belief is inferred from the value(s) of one or several other beliefs.
@@
We specify a prompt template for instantiating a prompt based on the agent’s beliefs, considering the agent’s name (the name of the person the agent represents), the name of the teacher, and information about why previous excuse attempts (if any) have failed. {minted}[fontsize=]js const genPrompt = beliefs =¿ ‘Can you write a charming yet convincing excuse for a student who forgot their homework? The names of teacher and student are ${beliefs.teacherName},and$beliefs.name, respectively (i.e., sign the excuse with ${beliefs.name}).\par Considerthefollowingfeedbackreceivedfrompastrejectedexcuses:\par$beliefs.rejectExps.map(exp =¿ ‘• ${exp}`).join(^{\prime}\n^{\prime})$‘ The prompt template is called in the body of the agent’s single plan. The body is activated (by the plan’s head) only if no excuse has been accepted so far. The body then constructs the excuse prompt and sends it to the LLM. When the agent has received the generated excuse, it returns the excuse to the environment. {minted}[fontsize=]js Plan( beliefs =¿ !beliefs.excuseAccepted, async beliefs =¿ const prompt = genPrompt(beliefs) const excusePromise = await model.generateContent([prompt]) const excuse = excusePromise.response.text() return excuse ) Above, we have instantiated a JS-son Plan object; later, the agent will apply a default deliberation function, thus saving us some implementation overhead.
@@
Next, we implement the agent’s belief revision function. JS-son’s default revision function simply merges percepts into the agent’s belief base, with percepts taking precedence over beliefs in case of conflicts. However, for our agent we need a more specific belief revision function that i) adds new rejection explanations (if novel) to our array of explanations and ii) updates the excuseAccepted belief based on the percepts. {minted}[fontsize=]js const reviseBeliefs = (beliefs, percepts) =¿ const rejectExps = percepts.rejectExp & !beliefs.rejectExps.includes(percepts.rejectExp) ? Array(…beliefs.rejectExps, percepts.rejectExp) : beliefs.rejectExps return percepts.excuseAccepted ? …beliefs, excuseAccepted: true, : …beliefs, rejectExps With our beliefs, plan, and belief revision function, we can then instantiate a JS-son Agent object. {minted}[fontsize=]js const agent = new Agent( id: ’student’, beliefs, [plan], reviseBeliefs ) Note that the agent’s runner function (next in JS-son terms) is built in, i.e., we do not need to implement it explicitly.
@@
As a prerequisite to implementing the environment, we need to specify the updateState function, of which we merely provide a sketch, abstracting from the exact checks executed when determining whether an excuse is acceptable or not. {minted}[fontsize=]js const updateState = async(actions, _, currentState) =¿ const excuse = await actions[0] if(!excuse) return …currentState, rejectExp: ” const state = excuseAccepted: false if (…) // apply rule to analyze excuse text state.rejectExp = … // explain rule violation else if (…) // apply another rule to analyze excuse text state.rejectExp = … // explain another rule violation else state.excuseAccepted = true return state As we can see, the function essentially applies a set of rules to assess whether the provided excuse is acceptable or not and, in case of a rule violation, provides an explanation for rejection (for the first rule that is violated, which terminates rule execution). Then, the state of the environment is updated accordingly (which implicitly leads to a corresponding update of the agent’s percepts).
@@
{minted}
[fontsize=]js const environment = new Environment( [agent], excuseAccepted: false, rejectExp: ” , updateState ) setInterval(() =¿ environment.run(1), 3000) Again, we do not explicitly implement a runner function, as this is provided by default by JS-son’s environment object. We run the environment in timed intervals. This reduces load on the LLM service API, but also smoothens the execution of a JS-son reasoning loop with asynchronous behavior—in contrast to asynchronous environments, asynchronous behavior within reasoning loops is not supported in a straight-forward, stable manner by JS-son at the time of writing  [^6] Cf. Appendix Item 4..
@@ 5.1 Centralized Multi-Agent Systems
We can implement the agent behavior we have sketched above as a single plan. {minted}[fontsize=]js Plan( () =¿ true, beliefs =¿ const neighborActivity = determineNeighborActivity( beliefs.index, beliefs.activityArray ) const isActive = beliefs.activityArray[beliefs.index] return ( isActive & neighborActivity ¿= 2 neighborActivity ¡ 4 ) —— neighborActivity === 3 ? nextRound: ’active’ : nextRound: ’inActive’ ) Note that the plan is always active (its head necessarily evaluates to true). The plan’s body depends on the following beliefs:
@@
Given our plan, we can generate the entire grid of agents by using a simple function that assigns the agents’ initial beliefs based on an initialActivity array that specifies the initial state of the agents (for example, in order to guarantee the emergence of a specific pattern). {minted}[fontsize=]js const generateAgents = initialActivity =¿ initialActivity.map((value, index) =¿ const beliefs = …Belief(’index’, index), …Belief(’activityArray’, initialActivity) return new Agent(index, beliefs, , [plan]) ) Next, we need to specify the environment’s behavior. Here, it is crucial to assure that updates are propagated at the end of a given round (cf. Figure 2a). Accordingly, our environment distinguishes between previous and next agent activity statuses, right from the start when specifying the initial state. {minted}[fontsize=]js const generateState = initialActivity =¿ ( previousActivity: initialActivity, nextActivity: [] ) Accordingly, the environment’s state update is propagated only when the last agent in a grid has submitted its action. {minted}[fontsize=]js const updateState = (actions, agentId, currentState) =¿ const stateUpdate = …currentState const agentActive = actions.some(action =¿ action.nextRound === ’active’ ) stateUpdate.nextActivity.push(agentActive) if (agentId == currentState.previousActivity.length - 1) return previousActivity: stateUpdate.nextActivity, nextActivity: [] return stateUpdate We also specify a state filter that exposes the agents only to the grid’s “previous” status, i.e., not to the status that is currently undergoing an update. {minted}[fontsize=]js const stateFilter = state =¿ ( activityArray: state.previousActivity ) Finally, we can generate the GoL’s initial activity grid and instantiate the environment.
{minted}
[fontsize=]js const initialActivity = generateInitialActivity() const env = new Environment( generateAgents(initialActivity), generateState(initialActivity), updateState, render, stateFilter ) With this we have finalized the specification of the GoL. We can, for example, attach the MAS to a rendering function that produces/manipulates HTML elements and run it in a Web browser [^7] Cf. Appendix Item 5..
@@ 5.2 Distributed Multi-Agent Systems
Let us start with the former. For this, we implement a higher-order function that generates our custom runner function that we will give to our shadow agent. The function the generator returns will send the belief update to the remote agent (instead of applying it to a locally running instance). {minted}[fontsize=]js const nextGen = agentId =¿ function next (beliefs) ws.send(JSON.stringify(beliefs)) return global.actionRequests[agentId] Now, we can instantiate the shadow agent, using JS-son’s RemoteAgent abstraction. {minted}[fontsize=]js const claustrophobe = new RemoteAgent( ’claustrophobe’, beliefs, nextGen(agentId) ) Let us move to environment-client communications. For this, we first implement a higher-order function generating a runner for the environment. Because we need the Web socket client to be available in the context of this function, we create the 3rd-order function runnerGen. The runner higher-order function provides a context in which the environment’s run function is executed. In our case, an execution cycle of the environment is triggered when a message of our remote agent arrives via the Web socket connection. {minted}[fontsize=]js const runnerGen = ws =¿ run =¿ () =¿ ws.on(’message’, message =¿ const jMessage = JSON.parse(message) if (jMessage.agentId & jMessage.actions) global.actionRequests[jMessage.agentId] = jMessage.actions run() ) On server side, we now have all components for implementing the environment (and running it). {minted}[fontsize=]js const environment = new Environment( [paranoid, claustrophobe, porter], state, update, state =¿ console.log(state), state =¿ state, runnerGen(ws) ) environment.run() Finally, we implement the client agent. Again, we abstract from agent-internal details, assuming we have a claustrophobe agent, whose reasoning loop we can execute.
To enable the Web socket server integration, we instantiate a client and send an initial opening request to the server to trigger the execution of the environment’s loop. {minted}[fontsize=]js ws.on(’open’, () =¿ ws.send(JSON.stringify( agentId, actions: [[]] )) ) Finally, we implement the Web socket listener that receives messages from the environment and triggers the agent’s reasoning loop using these messages as the percepts. {minted}[fontsize=]js ws.on(’message’, message =¿ const jMessage = JSON.parse(message) if (jMessage === Object(jMessage)) const actions = claustrophobe.next(jMessage) ws.send(JSON.stringify( agentId, actions )) ) We can then start server and client to run the MAS (leading to a loop in which the porter agent iteratively attempts to satisfy the wishes of the two other agents) [^8] Cf. Appendix Item 6..

### v2

@@
11institutetext: Department of Computing Science 22institutetext: Umeå University, 901 87 Umeå, Sweden, 22email: tkampik@cs.umu.se
Implementing Agents in JavaScript
[…]
@@ 4.1 A Simple Agent in Vanilla JavaScript
First, we model the beliefs of the agent.
const beliefs =
{ door: { locked: true },
requests: [] }
As we can see, the initial state of the belief base models that the door is currently locked, and that no requests are to be handled. Let us now specify the agent’s plans: i) if the door is not locked and a request to lock the door exists (head) then the agent should act to lock the door (body); ii) if the door is locked and a request to unlock the door exists (head) then the agent should act to unlock the door (body).
const plans = [
{ head:
beliefs =>
!beliefs.door.locked &&
beliefs.requests.includes(’lock’),
body: () => lock },
{ head: beliefs =>
beliefs.door.locked &&
beliefs.requests.includes(’unlock’),
body: () => unlock }
With beliefs and plans specified, we can move on to implement the agent’s reasoning loop. We start by implementing the belief revision function, which updates the agent’s belief about the state of the door, removes requests that have been successfully executed, and adds new requests that have come in.
const revise = (beliefs, percepts) => {
return {
door: percepts.door,
requests: beliefs.requests.filter(
request =>
!percepts.executions.includes(request)
).concat(percepts.requests)
Next, we implement the plan deliberation function. The function first determines which bodies of the plans should be activated (deliberate) to then execute these bodies (act).
const deliberate = (beliefs, plans) => {
return activatedBodies = plans.filter(
plan =>
plan.head(beliefs)
).map(plan => plan.body(beliefs))
Finally, we can complete the agent, by i) assigning beliefs and plans, as well as the revision and deliberation functions to an agent object, and ii) implementing a run function that triggers first belief revision and then plan deliberation.
const agent = {
beliefs,
plans,
revise,
deliberate,
get run() {
this.beliefs = this.revise(this.beliefs, percepts)
return this.deliberate(
this.beliefs,
this.plans
We can observe that we need to make use of a get function to access object-level properties and that percepts will need to be specified outside of the scope of the agent object or any of its functions. This indicates that providing generic abstractions in the form of a library may be useful.
We still need to implement the environment that provides our agent with percepts and processes its actions.
const environment = {
agent,
state: {
door: { locked: true },
requests: []
get run() {
[…Array(steps).keys()].forEach(
step => {
this.state.requests =
Math.random() < 0.5 ?
[’lock’] :
[’unlock’]
percepts = this.state
const exec = this.agent.run
const actions = exec ? exec : []
if (actions.includes(’lock’))
this.state.door.locked = true
if (actions.includes(’unlock’))
this.state.door.locked = false
this.state.executions = actions
@@
Our environment processes all agent actions by changing the state of the door accordingly and adds pseudo-random requests to either lock or unlock the door, independently of the door’s current status. In this way, the environment primitively simulates the presence of additional agents. Note that in this simple example, our environment is a single agent environment, i.e., it is not programmed to handle more than one agent. We again exploit the get syntax in a way that is arguably not idiomatic.
We can now execute the environment, first specifying the number of environment-agent interactions (steps) it should run.
const steps = 20
environment.run
@@
When running the example, we see that our porter agent processes the requests it receives from the environment as intended, generating the desired actions, which are then further processed accordingly by the environment [^5] Cf. Appendix Item 3.. When debugging nuances of the example, we observe that nonsensical requests (the locking of locked doors and the unlocking of unlocked doors) “pile up” until they can be meaningfully addressed.
@@ 4.2 An LLM Agent Based on Reusable Abstractions
We first implement the agent’s beliefs, by instantiating JS-son Belief objects.
const beliefs = {
Belief(’rejectExps’, []),
Belief(’excuseAccepted’, false),
Belief(’name’, Bart’),
Belief(’teacherName’, Edna Krabappel’)
In our case, the result is roughly the same as when specifying beliefs as JavaScript objects. However, in advanced scenarios, JS-son beliefs have in-built support for specifying revision behavior, such as priorities and inter-belief dependencies, where the value of one belief is inferred from the value(s) of one or several other beliefs.
@@
We specify a prompt template for instantiating a prompt based on the agent’s beliefs, considering the agent’s name (the name of the person the agent represents), the name of the teacher, and information about why previous excuse attempts (if any) have failed.
const genPrompt = beliefs =>
Can you write a charming yet convincing excuse
for a student who forgot their homework?
The names of teacher and student are
\${beliefs.teacherName}, and \${beliefs.name},
respectively (i.e., sign the excuse with \${beliefs.name}).
[…]
received from past rejected excuses:
[…]
The prompt template is called in the body of the agent’s single plan. The body is activated (by the plan’s head) only if no excuse has been accepted so far. The body then constructs the excuse prompt and sends it to the LLM. When the agent has received the generated excuse, it returns the excuse to the environment.
Plan(
beliefs => !beliefs.excuseAccepted,
async beliefs => {
const prompt = genPrompt(beliefs)
const excusePromise =
await model.generateContent([prompt])
const excuse = excusePromise.response.text()
return excuse
Above, we have instantiated a JS-son Plan object; later, the agent will apply a default deliberation function, thus saving us some implementation overhead.
@@
Next, we implement the agent’s belief revision function. JS-son’s default revision function simply merges percepts into the agent’s belief base, with percepts taking precedence over beliefs in case of conflicts. However, for our agent we need a more specific belief revision function that i) adds new rejection explanations (if novel) to our array of explanations and ii) updates the excuseAccepted belief based on the percepts.
const reviseBeliefs = (beliefs, percepts) => {
const rejectExps =
percepts.rejectExp &&
!beliefs.rejectExps.includes(percepts.rejectExp) ?
Array(…beliefs.rejectExps, percepts.rejectExp) :
beliefs.rejectExps
return percepts.excuseAccepted ? {
beliefs,
excuseAccepted: true,
beliefs,
rejectExps
With our beliefs, plan, and belief revision function, we can then instantiate a JS-son Agent object.
const agent = new Agent({ id: student’, beliefs, [plan], reviseBeliefs })
Note that the agent’s runner function (next in JS-son terms) is built in, i.e., we do not need to implement it explicitly.
As a prerequisite to implementing the environment, we need to specify the updateState function, of which we merely provide a sketch, abstracting from the exact checks executed when determining whether an excuse is acceptable or not.
const updateState = async(actions, _, currentState) => {
const excuse = await actions[0]
if(!excuse)
return { currentState, rejectExp: ”}
const state = {
excuseAccepted: false
if (…) { // apply rule to analyze excuse text
state.rejectExp = // explain rule violation
} else if (…) { // apply another rule to analyze excuse text
state.rejectExp = // explain another rule violation
} else {
state.excuseAccepted = true
return state
@@
As we can see, the function essentially applies a set of rules to assess whether the provided excuse is acceptable or not and, in case of a rule violation, provides an explanation for rejection (for the first rule that is violated, which terminates rule execution). Then, the state of the environment is updated accordingly (which implicitly leads to a corresponding update of the agent’s percepts).
@@
const environment = new Environment(
[agent],
{ excuseAccepted: false,
rejectExp: },
updateState
setInterval(() => environment.run(1), 3000)
Again, we do not explicitly implement a runner function, as this is provided by default by JS-son’s environment object. We run the environment in timed intervals. This reduces load on the LLM service API, but also smoothens the execution of a JS-son reasoning loop with asynchronous behavior—in contrast to asynchronous environments, asynchronous behavior within reasoning loops is not supported in a straight-forward, stable manner by JS-son at the time of writing  [^6] Cf. Appendix Item 4..
@@ 5.1 Centralized Multi-Agent Systems
We can implement the agent behavior we have sketched above as a single plan.
Plan(
() => true,
beliefs => {
const neighborActivity =
determineNeighborActivity(
beliefs.index, beliefs.activityArray
const isActive =
beliefs.activityArray[beliefs.index]
return (
isActive &&
neighborActivity >= 2 &&
neighborActivity < 4
) $|$ neighborActivity === 3 ?
{ nextRound: active } :
{ nextRound: inActive }
Note that the plan is always active (its head necessarily evaluates to true).
The plan’s body depends on the following beliefs:
@@
Given our plan, we can generate the entire grid of agents by using a simple function that assigns the agents’ initial beliefs based on an initialActivity array that specifies the initial state of the agents (for example, in order to guarantee the emergence of a specific pattern).
const generateAgents = initialActivity =>
initialActivity.map((value, index) => {
const beliefs = {
Belief(’index’, index),
Belief(’activityArray’, initialActivity)
return new Agent(index, beliefs, {}, [plan])
Next, we need to specify the environment’s behavior. Here, it is crucial to assure that updates are propagated at the end of a given round (cf. Figure 2a). Accordingly, our environment distinguishes between previous and next agent activity statuses, right from the start when specifying the initial state.
const generateState = initialActivity => ({
previousActivity: initialActivity,
nextActivity: []
Accordingly, the environment’s state update is propagated only when the last agent in a grid has submitted its action.
const updateState = (actions, agentId, currentState) => {
const stateUpdate = { currentState }
const agentActive =
actions.some(action =>
action.nextRound === active
stateUpdate.nextActivity.push(agentActive)
if (agentId == currentState.previousActivity.length - 1) {
return {
previousActivity: stateUpdate.nextActivity,
nextActivity: []
return stateUpdate
We also specify a state filter that exposes the agents only to the grid’s “previous” status, i.e., not to the status that is currently undergoing an update.
const stateFilter = state => (
{ activityArray: state.previousActivity }
Finally, we can generate the GoL’s initial activity grid and instantiate the environment.
const initialActivity = generateInitialActivity()
const env = new Environment(
generateAgents(initialActivity),
generateState(initialActivity),
updateState,
render,
stateFilter
With this we have finalized the specification of the GoL. We can, for example, attach the MAS to a rendering function that produces/manipulates HTML elements and run it in a Web browser [^7] Cf. Appendix Item 5..
@@ 5.2 Distributed Multi-Agent Systems
Let us start with the former. For this, we implement a higher-order function that generates our custom runner function that we will give to our shadow agent. The function the generator returns will send the belief update to the remote agent (instead of applying it to a locally running instance).
const nextGen = agentId => function next (beliefs) {
ws.send(JSON.stringify(beliefs))
return global.actionRequests[agentId]
Now, we can instantiate the shadow agent, using JS-son’s RemoteAgent abstraction.
const claustrophobe = new RemoteAgent(
claustrophobe’,
beliefs,
nextGen(agentId)
Let us move to environment-client communications. For this, we first implement a higher-order function generating a runner for the environment. Because we need the Web socket client to be available in the context of this function, we create the 3rd-order function runnerGen. The runner higher-order function provides a context in which the environment’s run function is executed. In our case, an execution cycle of the environment is triggered when a message of our remote agent arrives via the Web socket connection.
const runnerGen = ws => run => () => {
ws.on(’message’, message => {
const jMessage = JSON.parse(message)
if (jMessage.agentId && jMessage.actions)
global.actionRequests[jMessage.agentId] = jMessage.actions
run()
On server side, we now have all components for implementing the environment (and running it).
const environment = new Environment(
[paranoid, claustrophobe, porter],
state,
update,
state => console.log(state),
state => state,
runnerGen(ws)
environment.run()
Finally, we implement the client agent. Again, we abstract from agent-internal details, assuming we have a claustrophobe agent, whose reasoning loop we can execute.
To enable the Web socket server integration, we instantiate a client and send an initial opening request to the server to trigger the execution of the environment’s loop.
ws.on(’open’, () => {
ws.send(JSON.stringify({ agentId, actions: $[$]] }))
Finally, we implement the Web socket listener that receives messages from the environment and triggers the agent’s reasoning loop using these messages as the percepts.
ws.on(’message’, message => {
const jMessage = JSON.parse(message)
if (jMessage === Object(jMessage)) {
const actions = claustrophobe.next(jMessage)
ws.send(JSON.stringify({ agentId, actions }))
We can then start server and client to run the MAS (leading to a loop in which the porter agent iteratively attempts to satisfy the wishes of the two other agents) [^8] Cf. Appendix Item 6..